On January 1, 2017, the Sheridan Company ledger shows Equipment $48.400 and Accumulated Depreciation $17.960


1. On January 1, 2017, the Sheridan Company ledger shows Equipment $48.400 and Accumulated Depreciation $17.960. The depreciation resulted from using the straight-line method with a useful life of 10 years and a salvage value of $3,500. On this date, the company concludes that the equipment has a remaining useful life of only 2 years with the same salvage value. 


Compute the revised annual depreciation. 

The revised annual depreciation $_______ 


2. Suppose in its 2017 annual report that McDonald's Corporation reports beginning total assets of $29.80 billion, ending total assets of $31.60 billion, net sales of $21.85 billion, and net income of $4.20 billion

 (a) Compute McDonald's return on assets. (Round return on assets to 2 decimal places, e.g. 5.12%.) 

McDonald's return on assets = _______ %

 (b) Compute McDonald's asset turnover (Round asset turnover to 2 decimal places, e.g. 5.12.) 

McDonald's asset turnover times = _______ %
Answer #1

1) Revised annual depreciation = (48400-17960-3500)/2 = 13470 per year

2) Return on assets = Net income/average assets = 4.2/30.70 = 13.68%

Turnover = Sales/Average assets = 21.85/30.70 = 0.71 Times